OpenStudy (anonymous):

(2+x)/3x=1/4 find x (i know its -8 but i dont know how to get there)

OpenStudy (anonymous):

2+x= 3x/4 multi 4 8+4x= 3x sub 4x 8= -x -8=x

OpenStudy (anonymous):

thks